OpenStudy (anonymous):

For the function f(x) = x + 5, what is the ordered pair for the point on the graph when x = 4w? (4w, x + 5) (4w, 4w + 5) (x, x + 5) (x, 4w + 5)

OpenStudy (anonymous):

I would like a answer and a explanation so i can answer real ones better. (Example question)

OpenStudy (mathstudent55):

You are given the function f(x) = x + 5. To find f(any value) just plug in "any value" in for x. Examples: For x = 2, f(2) = 2 + 5 = 7; then the ordered pair is (2, 7) For x = 6, f(6) = 6 + 5 = 8; then the ordered pair is (6, 8) For x = p, f(p) = p + 5; then the ordered pair is (p, p + 5) For x = 3p, f(3p) = 3p + 5; then the ordered pair is (3p, 3p + 5) Now try your problem for x = 4w.